OpenStudy (anonymous):

There are five more children in Miss Brown's class than in Mr whites class. Ten children move from MrS.Browns class to mr. Whites class. now there are twice as many children in Mr. Whites as in Mrs. Browns room. How many children were in Miss browns room before the move?

OpenStudy (anonymous):

let the students be x and y respectively in miss B n Mr W class x-5=y x-10=2y

OpenStudy (anonymous):

aww sec one is wrong

OpenStudy (anonymous):

x-10=2(y+10)

OpenStudy (anonymous):

still its wrong